OpenStudy (anonymous):

How many terms are in the arithemetic sequence 2, 5, 8, ...., 293? recall An=A1+(n-1) d n and 1 are small

OpenStudy (anonymous):

im coming up with 98

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow 293 = 2 + 3(n - 1) }\) \(\Large \color{purple}{\rightarrow 293 = 2 + 3n - 3 }\) \(\Large \color{purple}{\rightarrow 293 = -1 + 3n }\) \(\Large \color{purple}{\rightarrow 294 = 3n }\) \(\Large \color{purple}{\rightarrow n = 98 }\)

Parth (parthkohli):

Yes, 293 is the 98th term.

OpenStudy (anonymous):

thank you once again parthkohli